OpenStudy (anonymous):

What is lim (1/b-1/2)/(b-2)? b->2

OpenStudy (anonymous):

3 squared

OpenStudy (anonymous):

how?

OpenStudy (123456789):

-(1/4) \[\lim_{b \rightarrow 2}\left( 1/b - 1/2 \right)/\left( b-2 \right)=\] \[\lim_{b \rightarrow 2}(\left( 2-b \right)/\left( 2b \right))/\left[ b-2 \right]\] \[\lim_{b \rightarrow 2}\left( 2-b \right)/\left( 2b*\left( b-2 \right) \right)\] \[\lim_{b \rightarrow 2} -1/(2b) =-1/4\]

OpenStudy (123456789):

did that help you?

OpenStudy (anonymous):

Thanks so much!